OpenStudy (maths99):

A trolley T moves at speed 1.2 m s–1 along a horizontal frictionless surface. The trolley collides with a stationary block on the end of a fixed spring. The mass of T is 250 g. T compresses the spring then expands and causes T to move in a direction opposite to its initial direction. At the time that T loses contact with the block, it is moving at a speed of 0.75 m s–1. Why is the change in momentum= m(v+u) and not m(v-u) since change in momentum is final momentum - initial.

OpenStudy (irishboy123):

\(\large u_i = 1.2 m/s, u_f = \color{red}{-0.75}m/s \) \(\large \Delta u = 1.2 - (-0.75)\) .....

OpenStudy (maths99):

Then shouldnt it be -0.75 - ( 1.2) @irishboy123

OpenStudy (irishboy123):

you get the same magnitude, just a different sign so yes, that's ok too

OpenStudy (maths99):

Thanks @irishboy123 really helped me alot

OpenStudy (irishboy123):

great!